OpenStudy (anonymous):

integrate sqrt(3x+1)/(3x+4)

hartnn (hartnn):

is everything under sqrt or just the numerator ?

OpenStudy (anonymous):

If it's the latter, I think a substitution would work.

OpenStudy (anonymous):

just numerator. Trying sub. getting messy.

hartnn (hartnn):

what substitution are you trying ?

OpenStudy (anonymous):

u=3x, du=3dx, 1/3du then split those integrals up after the sub. then sub again for t=sqrt(u), dt=2/3u^(3/2)

hartnn (hartnn):

no no

hartnn (hartnn):

try \(\large u=\sqrt{3x+1}\)

hartnn (hartnn):

denominator = u^2+3

hartnn (hartnn):

du = .. ?

OpenStudy (anonymous):

1/2(3x+1)^-1/2*3

hartnn (hartnn):

or can i say , du = 3/ (2u) dx ?

hartnn (hartnn):

and substitute (2u du)/3 in place of dx

hartnn (hartnn):

getting this ?

OpenStudy (anonymous):

yeah never done or thought about it like that.

hartnn (hartnn):

time to try out new things! :) you will do this again in many problems, for sure.

OpenStudy (anonymous):

thanks

hartnn (hartnn):

oh, so will you be able to solve this now ?

OpenStudy (anonymous):

I think so, the 3x+4 on the denominator stays though when i sub right.

hartnn (hartnn):

3x+4 = 3x+1 + 3 = u^2+3 we need to get rid of every x when we substitute...

OpenStudy (anonymous):

right just confused becasue u was sqrt(3x+1) not just 3x+1

hartnn (hartnn):

yeah..